PartnerObserving Log for 2007 Oct 04-05

Observers: J. Hill, V. Testa
Instrument Support: R. Bertram
Telescope Operator: A. Ceranski

This is scheduled SDT time which we have commandeered to finish the technical observing engineering activities which are needed for the SDT observing. (5.5 of the last 6 technical observing nights have been weathered out.)

00:11 - Five biases

Open at 01:10 UT. T=9degC D=7degC Wind=9m/s from south.

Skyflats

Vincenzo took a set of Uspec and V skyflats in non-photometric sky pointed at 18:30 +17, EL=70.

01:26 - SkyFlatTest Uspec sequence, 1sec =8700 counts

01:30 - Start of the Uspec SkyFlats at PA=0, 7 sec [24000,............16000]

01:35 - SkyFlatTest V sequence, 1 sec = 5000 counts

01:37 - Start of the V SkyFlats at PA=0 , 12 sec [35000,......]]

Zeropoint of the pointing model

We are adjusting the zeropoint of the pointing model using a star on the meridian in the south = 19:06:14.94 -04:52:57.1 . Editting the corrections into the file /usr/commonsw/Configuration/PCS/LBCLEFTpointingmodel.dat (so far on lbtmu102 only).

Nominal Zeropoint IA=-289.7 IE=-57.9 arcsec
015041 - V 1sec, star at X=805 Y=3089
Next ==> dIA=+89 arcsec
015458 - V1sec, star at X=561 Y=3061 (wrong way)
Next ==> dIA==180 arcsec
015812 - V1sec, star at X=1041 Y=3125 (OK in Azimuth)
Next ==> dIE=-40 arcsec
020206 - V1sec, Clouds
020454 - V5sec, Clouds
020935 - V50sec, star at X=1069 Y=2965 (near the nominal rotator center)
Corrected Zeropoint IA=-380 IE=-98 arcsec

02:13 - Close for clouds and rising humidity

02:15 - Five biases

Dark tests with the LBC-Blue front cover on

07:16 - 900 sec dark with shutter closed (shows flat-field-like pattern).

07:43 - 900 sec V exposure with shutter open and cover on. Image 074316 was lost to Trackers CCD controller error. See Issue #839.

08:29 - 25 biases after power cycling the camera

08:25 - 900 sec V exposure

08:49 - CCD controller problem happens again; power cycle again

092535 - 900 sec dark with shutter closed

10:09 - Trackers CCD controller error again; power cycle again.

10:15 - Spent next hour fooling around with some kind of PCS/TCS hang up. Wound up having to restart all of TCS. See Issues #840 and #841. 840 was probably a precursor to 841. Fog continues.

10:41 - lbckill/lbcstart, then Trackers controller error again at 10:50:25

10:53 - reboot CMU

11:18 - 5 biases

112228 - 900 sec V exposure with shutter open and telescope tracking at EL=16, ROT=0

114115 - 900 sec dark exposure with shutter closed, ROT=0

Results from these exposures are that we see 4 counts/min (60 counts in 900 sec) of excess light above dark current with the LBC-Blue front cover on. Combined with last night's data, this means that 60% of the dome-closed stray light is coming from outside the hub, and 40% from inside the hub. Aaargh.

11:53:53 UI freezes on lbtmu03, but not on lbtmu05. Unfreezes by itself at 11:56:05 (8 sec before exposure ends).

Night ends with no break in the humidity. Forecast is better for tomorrow.

More CCD tests in the afternoon

21:12 - Turn on; get Profibus error on rotator; Turn off

21:16 - Turn on
212500 - 50 sec U exposure
212825 - 50 sec U exposure
21:57 - first of a sequence of 50 sec U exposures
22:15:50 - RPC Communication error from Science controller; after 8 exposures; Turn off

22:32 - Turn on
223653 - first of a sequence of 50 sec U exposures
23:04:46 - RPC Communication error from Trackers controller
23:38 - Turn off and reboot CMU

23:45 - Turn on
234936 - first of a sequence of 50 sec U exposures (ran until Vincenzo stopped them to take biases)

So, good old traditional LBC errors have appeared, but we see no sign of the Bias Voltage errors or the Trackers CCD controller errors that have been bothering us for the last two days? Dewpoint has dropped from 7 degC, to 4degC. Dome temperature still about 10degC for these tests. ???

-- JohnHill - 05 Oct 2007
Topic revision: r5 - 21 Jul 2009, NormCushing
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback